OpenStudy (griffindore):

Which expression is equivalent to 2^4 ⋅ 2^−7 1/2^11 1/2^3 2^3 2^11

Nnesha (nnesha):

when you multiply same bases you should add their exponents \[\huge\rm x^m \times x^n =x^{m+n}\]

OpenStudy (griffindore):

oh s0 2^11?

Nnesha (nnesha):

nope

Nnesha (nnesha):

2nd exponent is negative so positive - negative = ?

OpenStudy (griffindore):

i meant -11

OpenStudy (griffindore):

oh minus

OpenStudy (griffindore):

so that makes it c right?

Nnesha (nnesha):

no here is an example \[\huge\rm x^m \times x^{-n} = x^{m+(-n)}\]

OpenStudy (griffindore):

so it b??? im a litlle confused

Nnesha (nnesha):

explain why do you think its b ? don't guess

OpenStudy (griffindore):

because of the negative and there is no other awnser that maes sence so proccess of eliminaion

Nnesha (nnesha):

what exponent did you get at the top ?

OpenStudy (griffindore):

-3

Nnesha (nnesha):

yep so you are not allowed to leave negative exponent so flip the fraction

Nnesha (nnesha):

so your answer is correct.

OpenStudy (griffindore):

thanks

Nnesha (nnesha):

np